# Reproducing Kernel Hilbert Spaces on Banach Completions of Virtual Persistence Diagram Groups

### v1

@@ 1 Introduction
Persistent homology assigns to a filtered simplicial complex a persistence diagram: a finite multiset of points in $\mathbb{R}^{2}$ encoding the birth and death parameters of homological features. Under the usual tameness assumptions ensuring finitely many homological critical values, the resulting diagram is Lipschitz-stable on the input. More precisely, the map from filtrations (or functions) equipped with the $\|\cdot\|_{\infty}$ metric to persistence diagrams equipped with the bottleneck distance is 11–Lipschitz, and, under appropriate summability conditions on lifetimes, is also Lipschitz with respect to pp–Wasserstein distances [892133, CohenSteiner2007StabilityPD, Oudot2015PersistenceT].
@@
Kernel and embedding methods for persistence diagrams are by now well developed. Prominent examples include persistence images, which map diagrams to finite-dimensional $L^{2}$-type representations [adams2017persistenceimages], the persistence scale-space kernel [reininghaus2014stablemultiscale], the persistence weighted Gaussian kernel [kusano2016pwgk], the Sliced Wasserstein kernel [carriere2017slicedwasserstein], and the persistence Fisher kernel [le2018persistencefisher]. However, they do not operate on the intrinsic algebraic structure of the space of persistence diagrams itself. That space is a free commutative, cancellative monoid generated by the birth–death space, with monoid operation given by addition of multiplicities and equipped with bottleneck or Wasserstein matching metrics. Since ordinary persistence diagrams do not form an LCA group, these constructions do not have access to the canonical functorial identification as in the case of Pontryagin duality and Bochner’s theorem.
@@
Virtual persistence diagrams, introduced by Bubenik and Elchesen [bubenik2022virtualpd], are obtained by adjoining additive inverses to finite persistence diagrams in a manner compatible with matching metrics. Given a metric pair $(X,d,A)$, let $D(X,A)$ denote the commutative monoid of finite diagrams relative to $A$, and let $K(X,A)$ denote its Grothendieck group. For the Wasserstein–$1$ distance, [bubenik2022virtualpd] defines a translation-invariant metric $\rho$ on $K(X,A)$ extending the Wasserstein–$1$ metric on $D(X,A)$. More generally, for $p\in[1,\infty]$, the $p$-Wasserstein distance on $D(X,A)$ is translation invariant if and only if the associated strengthened metric $d_{p}$ is a $p$-metric [bubenik2022virtualpd]. In particular, $p=1$ is the unique exponent for which translation invariance holds for every metric pair. Since translation invariance is required to extend a matching metric from diagrams to the Grothendieck group, this identifies the Wasserstein–$1$ metric as the canonical choice for virtual persistence diagram groups.
@@
Beyond the group level, the Grothendieck construction admits a canonical linearization. The group $(K(X,A),\rho)$ embeds isometrically into a normed vector space $(V(X,A),\|\cdot\|_{W_{1}})$, and its Banach completion $B:=\widehat{V}(X,A)$ provides a canonical Banach-space model of the $W_{1}$ geometry of virtual persistence diagrams [bubenik2022virtualpd]. Under the standing assumptions fixed in Section 2 (in particular, that the quotient by $A$ equipped with the $1$-strengthened metric yields a separated pointed metric space), one has a canonical isometric isomorphism
@@
the Lipschitz-free (Arens–Eells) Banach space over the pointed metric space $(X/A,d_{1},[A])$ [bubenik2022virtualpd, weaver2018lipschitz].
In the uniformly discrete case, the metric group $(K(X,A),\rho)$ is discrete and locally compact, and the duality between $K(X,A)$ and its Pontryagin dual supports a correspondence in which analytic constructions on $K(X,A)$ can be recovered from their dual representations, as recalled in Section 2.2 [Folland2015CourseAH, fanning2025reproducingkernelhilbertspaces]. In Section 3, we show that this is the only situation in which such a correspondence is available for the metric topology induced by $\rho$: $(K(X,A),\rho)$ is locally compact if and only if it is discrete, and this holds exactly when the pointed metric space $(X/A,d_{1},[A])$ is uniformly discrete. The argument combines general properties of the Grothendieck metric with rigidity results for locally compact group topologies on free abelian groups, as stated, for example, in [morris1977locallycompact].
@@ 1.2 Related work
Foundational results on persistent homology, barcodes/persistence diagrams, and stability were established in [892133, Zomorodian2005ComputingPH, CohenSteiner2007StabilityPD, Oudot2015PersistenceT], with subsequent structural developments including stability and structure theorems for persistence modules [chazal2013structurestabilitypersistencemodules]. Extensions such as extended persistence [CohenSteiner2009ExtendingPU], zigzag persistence [carlsson2008zigzagpersistence], and vineyards [CohenSteiner2006VinesAV] broaden the classical one-parameter setting, while multiparameter persistence replaces complete discrete invariants by interleaving-based analysis [Lesnick2015InterleavingMP, chazal2013structurestabilitypersistencemodules]. Virtual persistence diagrams and their Grothendieck/Banach-space formalism are developed in [bubenik2022virtualpd]. RKHS constructions based on multipliers on the Pontryagin dual in the uniformly discrete (locally compact) regime are developed in [fanning2025reproducingkernelhilbertspaces].
@@ 2.1 Virtual persistence diagrams
This characterization is due to [bubenik2022virtualpd]. In particular, $W_{1}[d]$ is translation invariant for every metric pair, whereas for metric spaces with nontrivial length geometry, translation invariance fails for all $p>1$. Since translation invariance is necessary for extending the Wasserstein distance to the Grothendieck group of diagrams, the choice $p=1$ is forced and canonical.
@@
The function $\rho$ is a well-defined translation-invariant metric on $K(X,A)$ [bubenik2022virtualpd] and extends $W_{1}$ along the canonical embedding $D(X,A)\hookrightarrow K(X,A)$.
Let $V(X)$ denote the free real vector space on $X$ and define $V(X,A):=V(X)/V(A)$. [bubenik2022virtualpd] defines a canonical norm $\|\cdot\|_{W_{1}}$ on $V(X,A)$ extending $\rho$ on $K(X,A)$. Let $\widehat{V}(X,A)$ denote the Banach completion.
@@
There is a chain of isometric embeddings [bubenik2022virtualpd]
@@ 2.2 Reproducing Kernel Hilbert Spaces for Finite Virtual Persistence Diagrams
Following [fanning2025reproducingkernelhilbertspaces], the quotient metric space $(X/A,d_{1})$ is modeled by a finite connected weighted graph whose shortest-path metric agrees with $d_{1}$. The associated graph Laplacian induces a nonnegative Dirichlet symbol
@@
The main analytic results of [fanning2025reproducingkernelhilbertspaces] provide explicit global Lipschitz bounds for functions in $H_{t}$ with respect to the Grothendieck metric $\rho$. In spectral form, one obtains the following estimate.
Corollary 2.4 (Spectral form [fanning2025reproducingkernelhilbertspaces, Corollary 2]).
@@
Corollary 2.5 (Geometric form [fanning2025reproducingkernelhilbertspaces, Corollary 3]).
@@ 3 Classification of Virtual Persistence Diagrams
This appears as Theorem 31 in [morris1977locallycompact].
@@ 4 Main Results
By Corollary 3.3, whenever $(K(X,A),\rho)$ is non-discrete, the metric topology induced by $\rho$ is not locally compact. The results of this section therefore proceed by working with the Banach completion $B:=\widehat{V}(X,A)$ of $K(X,A)$ with respect to $\|\cdot\|_{W_{1}}$, which is canonically isometric to the Lipschitz-free (Arens–Eells) Banach space $\mathcal{F}(X/A,d_{1})$ by [bubenik2022virtualpd, Section 7].
@@ 4.2 Structural consequences
By [bubenik2022virtualpd, Corollary 7.10], the space $B$ is canonically isometric to the Lipschitz-free Banach space $\mathcal{F}(X/A,d_{1})$. Lipschitz-free Banach spaces over separable pointed metric spaces are separable; see [weaver2018lipschitz], Section 3.1. The claim follows directly from (H1). ∎
@@ 4.3 Gaussian kernels induced by Hilbert embeddings
A finite subset $S\subset X/A\setminus\{[A]\}$ together with a sign pattern $\sigma\colon S\to\{\pm 1\}$ is called basepoint-separated if $d_{1}(u,v)\geq d_{1}(u,[A])+d_{1}(v,[A])$ whenever $\sigma(u)\neq\sigma(v).$ For each basepoint-separated pair $(S,\sigma)$, define $v([A])=0$ and $v(u)=\sigma(u)\,d_{1}(u,[A])$ for $u\in S$, and let $f_{S,\sigma}(x):=\inf_{y\in S\cup\{[A]\}}\bigl(v(y)+d_{1}(x,y)\bigr)\in\mathrm{Lip}_{0}(X/A,d_{1})$ denote the associated McShane extension [weaver2018lipschitz]. Then $\mathrm{Lip}_{d_{1}}(f_{S,\sigma})\leq 1$ and $f_{S,\sigma}(u)=\sigma(u)\,d_{1}(u,[A])$ for all $u\in S$. We assume that whenever a basepoint-separated pair $(S,\sigma)$ is used below, the corresponding witness $f_{S,\sigma}$ occurs among the coordinates $(\ell_{n})_{n\geq 1}$. Fix a strictly positive weight sequence $(w_{n})_{n\geq 1}\in\ell^{2}$.
@@ 4.4 Comparison of covariance structures
To obtain the Rayleigh bounds [10.5555/280490], fix $x\in B$ with $\langle Q_{1}x,x\rangle>0$, equivalently $s_{1}(x)>0$. Dividing the two seminorm inequalities by $s_{1}(x)$ and squaring yields
@@
Taking the infimum and supremum over $\{x:\langle Q_{1}x,x\rangle>0\}$ gives exactly the displayed Rayleigh bounds [10.5555/280490]. This completes the proof. ∎
@@ 6 Conclusion
The central structural result of this paper is a complete classification of local compactness for the Grothendieck metric group $(K(X,A),\rho)$. Theorem 3.2 shows that $(K(X,A),\rho)$ is locally compact if and only if its metric topology is discrete, and Corollary 3.3 identifies this condition with uniform discreteness of the pointed metric space $(X/A,d_{1},[A])$. In the uniformly discrete case, $K(X,A)\cong\bigoplus_{x\in X\setminus A}\mathbb{Z}$ is a discrete locally compact abelian group, and the Fourier and RKHS constructions based on Pontryagin duality and Haar measure apply as recalled in Section 2.2. Outside this case, which is the generic situation under the standing hypotheses (H1)–(H2) in Definition 2, $(K(X,A),\rho)$ is non-locally compact and Bochner’s theorem is unavailable. To treat this non-discrete case, we pass to the Banach completion $B=\widehat{V}(X,A)$, which is canonically isometric to the Lipschitz-free space $\mathcal{F}(X/A,d_{1})$ and infinite dimensional by Lemmas 4.1 and 4.2. From a norming family $(\ell_{n})\subset B^{\ast}$, weights $(w_{n})\in\ell^{2}$, and a diagonal trace-class covariance operator $\Sigma$, Section 4.3 constructs a bounded linear map $J\colon B\to\ell^{2}$ and an associated centered Gaussian measure with covariance operator $Q_{J,\Sigma}$. Theorem 4.3 shows that this data determines, for each $t>0$, a translation-invariant positive definite kernel $k_{J,\Sigma,t}(x,y)=\exp\!\left(-\frac{t}{2}\,\|\Sigma^{1/2}J(x-y)\|_{\ell^{2}}^{2}\right).$ Section 4.4 shows that every continuous Hilbertian seminorm on $B$ is realized by such covariance data, and Theorem 4.7 gives an exact criterion, in terms of Rayleigh quotients [10.5555/280490] of the associated covariance operators, for when two choices induce uniformly equivalent Gaussian kernels. When these kernels are restricted to $K(X,A)$, Theorem 4.4 provides an explicit global bound on the $\rho$-Lipschitz constants of all functions in the corresponding reproducing kernel Hilbert spaces, while Theorem 4.6 bounds the covering numbers of subsets of $K(X,A)$ in the induced feature metric by those of the Grothendieck metric. In the opposite direction, Theorem 4.5 shows that a single kernel value $k_{J,\Sigma,t}(g,0)$, together with finite support information, yields an explicit upper bound on the diagrammatic mass $\mathcal{M}(g)$. Section 4.5 then shows that all of these bounds persist, up to explicit probabilistic error terms, when kernel evaluations are replaced by random Fourier feature approximations.
@@
  • Authors’ Contributions C.F. developed the theoretical framework, conducted the experiments, and wrote the manuscript. M.E.A. advised the project and provided feedback on the framework, experiments, and manuscript.

### v2

@@ 1 Introduction
Persistent homology assigns to a filtered simplicial complex a persistence diagram: a finite multiset of points in $\mathbb{R}^{2}$ encoding the birth and death parameters of homological features. Under the usual tameness assumptions ensuring finitely many homological critical values, the resulting diagram is Lipschitz-stable on the input. More precisely, the map from filtrations (or functions) equipped with the $\|\cdot\|_{\infty}$ metric to persistence diagrams equipped with the bottleneck distance is 11–Lipschitz, and, under appropriate summability conditions on lifetimes, is also Lipschitz with respect to pp–Wasserstein distances [1, 2, 3].
@@
Kernel and embedding methods for persistence diagrams are by now well developed. Prominent examples include persistence images, which map diagrams to finite-dimensional $L^{2}$-type representations [4], the persistence scale-space kernel [5], the persistence weighted Gaussian kernel [6], the Sliced Wasserstein kernel [7], and the persistence Fisher kernel [8]. However, they do not operate on the intrinsic algebraic structure of the space of persistence diagrams itself. That space is a free commutative, cancellative monoid generated by the birth–death space, with monoid operation given by addition of multiplicities and equipped with bottleneck or Wasserstein matching metrics. Since ordinary persistence diagrams do not form an LCA group, these constructions do not have access to the canonical functorial identification as in the case of Pontryagin duality and Bochner’s theorem.
@@
Virtual persistence diagrams, introduced by Bubenik and Elchesen [9], are obtained by adjoining additive inverses to finite persistence diagrams in a manner compatible with matching metrics. Given a metric pair $(X,d,A)$, let $D(X,A)$ denote the commutative monoid of finite diagrams relative to $A$, and let $K(X,A)$ denote its Grothendieck group. For the Wasserstein–$1$ distance, [9] defines a translation-invariant metric $\rho$ on $K(X,A)$ extending the Wasserstein–$1$ metric on $D(X,A)$. More generally, for $p\in[1,\infty]$, the $p$-Wasserstein distance on $D(X,A)$ is translation invariant if and only if the associated strengthened metric $d_{p}$ is a $p$-metric [9]. In particular, $p=1$ is the unique exponent for which translation invariance holds for every metric pair. Since translation invariance is required to extend a matching metric from diagrams to the Grothendieck group, this identifies the Wasserstein–$1$ metric as the canonical choice for virtual persistence diagram groups.
@@
Beyond the group level, the Grothendieck construction admits a canonical linearization. The group $(K(X,A),\rho)$ embeds isometrically into a normed vector space $(V(X,A),\|\cdot\|_{W_{1}})$, and its Banach completion $B:=\widehat{V}(X,A)$ provides a canonical Banach-space model of the $W_{1}$ geometry of virtual persistence diagrams [9]. Under the standing assumptions fixed in Section 2 (in particular, that the quotient by $A$ equipped with the $1$-strengthened metric yields a separated pointed metric space), one has a canonical isometric isomorphism
@@
the Lipschitz-free (Arens–Eells) Banach space over the pointed metric space $(X/A,d_{1},[A])$ [9, 10].
In the uniformly discrete case, the metric group $(K(X,A),\rho)$ is discrete and locally compact, and the duality between $K(X,A)$ and its Pontryagin dual supports a correspondence in which analytic constructions on $K(X,A)$ can be recovered from their dual representations, as recalled in Section 2.2 [11, 12]. In Section 3, we show that this is the only situation in which such a correspondence is available for the metric topology induced by $\rho$: $(K(X,A),\rho)$ is locally compact if and only if it is discrete, and this holds exactly when the pointed metric space $(X/A,d_{1},[A])$ is uniformly discrete. The argument combines general properties of the Grothendieck metric with rigidity results for locally compact group topologies on free abelian groups, as stated, for example, in [13].
@@ 1.2 Related work
Foundational results on persistent homology, barcodes/persistence diagrams, and stability were established in [1, 14, 2, 3], with subsequent structural developments including stability and structure theorems for persistence modules [15]. Extensions such as extended persistence [16], zigzag persistence [17], and vineyards [18] broaden the classical one-parameter setting, while multiparameter persistence replaces complete discrete invariants by interleaving-based analysis [19, 15]. Virtual persistence diagrams and their Grothendieck/Banach-space formalism are developed in [9]. RKHS constructions based on multipliers on the Pontryagin dual in the uniformly discrete (locally compact) regime are developed in [12].
@@ 2.1 Virtual persistence diagrams
This characterization is due to [9]. In particular, $W_{1}[d]$ is translation invariant for every metric pair, whereas for metric spaces with nontrivial length geometry, translation invariance fails for all $p>1$. Since translation invariance is necessary for extending the Wasserstein distance to the Grothendieck group of diagrams, the choice $p=1$ is forced and canonical.
@@
The function $\rho$ is a well-defined translation-invariant metric on $K(X,A)$ [9] and extends $W_{1}$ along the canonical embedding $D(X,A)\hookrightarrow K(X,A)$.
Let $V(X)$ denote the free real vector space on $X$ and define $V(X,A):=V(X)/V(A)$. [9] defines a canonical norm $\|\cdot\|_{W_{1}}$ on $V(X,A)$ extending $\rho$ on $K(X,A)$. Let $\widehat{V}(X,A)$ denote the Banach completion.
@@
There is a chain of isometric embeddings [9]
@@ 2.2 Reproducing Kernel Hilbert Spaces for Finite Virtual Persistence Diagrams
Following [12], the quotient metric space $(X/A,d_{1})$ is modeled by a finite connected weighted graph whose shortest-path metric agrees with $d_{1}$. The associated graph Laplacian induces a nonnegative Dirichlet symbol
@@
The main analytic results of [12] provide explicit global Lipschitz bounds for functions in $H_{t}$ with respect to the Grothendieck metric $\rho$. In spectral form, one obtains the following estimate.
Corollary 2.4 (Spectral form [12, Corollary 2]).
@@
Corollary 2.5 (Geometric form [12, Corollary 3]).
@@ 3 Classification of Virtual Persistence Diagrams
This appears as Theorem 31 in [13].
@@ 4 Main Results
By Corollary 3.3, whenever $(K(X,A),\rho)$ is non-discrete, the metric topology induced by $\rho$ is not locally compact. The results of this section therefore proceed by working with the Banach completion $B:=\widehat{V}(X,A)$ of $K(X,A)$ with respect to $\|\cdot\|_{W_{1}}$, which is canonically isometric to the Lipschitz-free (Arens–Eells) Banach space $\mathcal{F}(X/A,d_{1})$ by [9, Section 7].
@@ 4.2 Structural consequences
By [9, Corollary 7.10], the space $B$ is canonically isometric to the Lipschitz-free Banach space $\mathcal{F}(X/A,d_{1})$. Lipschitz-free Banach spaces over separable pointed metric spaces are separable; see [10], Section 3.1. The claim follows directly from (H1). ∎
@@ 4.3 Gaussian kernels induced by Hilbert embeddings
A finite subset $S\subset X/A\setminus\{[A]\}$ together with a sign pattern $\sigma\colon S\to\{\pm 1\}$ is called basepoint-separated if $d_{1}(u,v)\geq d_{1}(u,[A])+d_{1}(v,[A])$ whenever $\sigma(u)\neq\sigma(v).$ For each basepoint-separated pair $(S,\sigma)$, define $v([A])=0$ and $v(u)=\sigma(u)\,d_{1}(u,[A])$ for $u\in S$, and let $f_{S,\sigma}(x):=\inf_{y\in S\cup\{[A]\}}\bigl(v(y)+d_{1}(x,y)\bigr)\in\mathrm{Lip}_{0}(X/A,d_{1})$ denote the associated McShane extension [10]. Then $\mathrm{Lip}_{d_{1}}(f_{S,\sigma})\leq 1$ and $f_{S,\sigma}(u)=\sigma(u)\,d_{1}(u,[A])$ for all $u\in S$. We assume that whenever a basepoint-separated pair $(S,\sigma)$ is used below, the corresponding witness $f_{S,\sigma}$ occurs among the coordinates $(\ell_{n})_{n\geq 1}$. Fix a strictly positive weight sequence $(w_{n})_{n\geq 1}\in\ell^{2}$.
@@ 4.4 Comparison of covariance structures
To obtain the Rayleigh bounds [20], fix $x\in B$ with $\langle Q_{1}x,x\rangle>0$, equivalently $s_{1}(x)>0$. Dividing the two seminorm inequalities by $s_{1}(x)$ and squaring yields
@@
Taking the infimum and supremum over $\{x:\langle Q_{1}x,x\rangle>0\}$ gives exactly the displayed Rayleigh bounds [20]. This completes the proof. ∎
@@ 6 Conclusion
The central structural result of this paper is a complete classification of local compactness for the Grothendieck metric group $(K(X,A),\rho)$. Theorem 3.2 shows that $(K(X,A),\rho)$ is locally compact if and only if its metric topology is discrete, and Corollary 3.3 identifies this condition with uniform discreteness of the pointed metric space $(X/A,d_{1},[A])$. In the uniformly discrete case, $K(X,A)\cong\bigoplus_{x\in X\setminus A}\mathbb{Z}$ is a discrete locally compact abelian group, and the Fourier and RKHS constructions based on Pontryagin duality and Haar measure apply as recalled in Section 2.2. Outside this case, which is the generic situation under the standing hypotheses (H1)–(H2) in Definition 2, $(K(X,A),\rho)$ is non-locally compact and Bochner’s theorem is unavailable. To treat this non-discrete case, we pass to the Banach completion $B=\widehat{V}(X,A)$, which is canonically isometric to the Lipschitz-free space $\mathcal{F}(X/A,d_{1})$ and infinite dimensional by Lemmas 4.1 and 4.2. From a norming family $(\ell_{n})\subset B^{\ast}$, weights $(w_{n})\in\ell^{2}$, and a diagonal trace-class covariance operator $\Sigma$, Section 4.3 constructs a bounded linear map $J\colon B\to\ell^{2}$ and an associated centered Gaussian measure with covariance operator $Q_{J,\Sigma}$. Theorem 4.3 shows that this data determines, for each $t>0$, a translation-invariant positive definite kernel $k_{J,\Sigma,t}(x,y)=\exp\!\left(-\frac{t}{2}\,\|\Sigma^{1/2}J(x-y)\|_{\ell^{2}}^{2}\right).$ Section 4.4 shows that every continuous Hilbertian seminorm on $B$ is realized by such covariance data, and Theorem 4.7 gives an exact criterion, in terms of Rayleigh quotients [20] of the associated covariance operators, for when two choices induce uniformly equivalent Gaussian kernels. When these kernels are restricted to $K(X,A)$, Theorem 4.4 provides an explicit global bound on the $\rho$-Lipschitz constants of all functions in the corresponding reproducing kernel Hilbert spaces, while Theorem 4.6 bounds the covering numbers of subsets of $K(X,A)$ in the induced feature metric by those of the Grothendieck metric. In the opposite direction, Theorem 4.5 shows that a single kernel value $k_{J,\Sigma,t}(g,0)$, together with finite support information, yields an explicit upper bound on the diagrammatic mass $\mathcal{M}(g)$. Section 4.5 then shows that all of these bounds persist, up to explicit probabilistic error terms, when kernel evaluations are replaced by random Fourier feature approximations.
@@
  • Authors’ Contributions C.F. developed the theoretical framework, conducted and analyzed the examples, and wrote the manuscript. M.E.A. advised the project and provided feedback on the framework and manuscript.